# Retained-spin micropolar hydrodynamics from the Boltzmann–Curtiss equation: a generalized Chapman–Enskog construction

### 31 Mar 2026 v1

@@ Abstract
We derive a retained-spin micropolar hydrodynamic closure from the Boltzmann–Curtiss equation using a generalized Chapman–Enskog construction in which the local mean spin is retained as a quasi-slow variable. Starting from the exact kinetic balance laws for mass, linear momentum, and intrinsic angular momentum, we isolate the bookkeeping relation between antisymmetric stress and stress-induced spin torque, decompose the first-order source into irreducible scalar, axial, and symmetric-traceless sectors, and show explicitly how the standard micropolar constitutive structure with coefficients $(\eta,\xi,\eta_{r},\alpha,\beta,\gamma)$ emerges. This decomposition makes clear that the one-particle kinetic stress contributes only to the symmetric stress, whereas the rotational viscosity belongs to an intrinsic/collisional transfer channel. For perfectly rough elastic hard spheres, we further obtain explicit dilute-gas estimates for the rotational viscosity $\eta_{r}$ from homogeneous spin relaxation and for the transverse spin-diffusion combination $\beta+\gamma$ from a transport-relaxation calculation. Targeted event-driven molecular-dynamics simulations are used as a posteriori checks: expanded homogeneous-spin density and roughness sweeps support the predicted $n^{2}$ and $K/(K+1)$ trends for $\eta_{r}$, while finite-$k$ transverse runs provide a qualitative diagnostic of the retained-spin response. The result is a self-contained derivation and coefficient-level estimate of retained-spin micropolar hydrodynamics that clarifies which parts of the closure are exact, which are first-order generalized Chapman–Enskog results, and which remain controlled rough-sphere estimates.
@@ I Introduction
The basic links between the kinetic and continuum descriptions are classical, but in practice the derivation is scattered across several literatures. The exact moment balances are often stated without algebra. The Chapman–Enskog construction is typically described only in outline. The antisymmetric stress channel, which is the key to rotational viscosity, is especially susceptible to confusion because it is not generated by the symmetric one-particle kinetic stress. In addition, when a local spin field is retained at the hydrodynamic level, the calculation is no longer a strict hydrodynamic reduction in the narrow sense: the mean spin is not a collision invariant in general and must be treated as a quasi-slow variable in the sense of extended thermodynamics [19, 13].
@@
Subsequent developments also proceeded largely along separate tracks. Micropolar continuum theory matured as an independent constitutive framework [10, 11]; granular and rough-sphere kinetic theories emphasized rapid granular flow, rough-particle transport, and hydrodynamic closures for particles with rotational degrees of freedom [14, 12, 17]; and recent Boltzmann–Curtiss-based studies have focused either on local-spin first-order closures or on higher-order constitutive models for polyatomic gases [26, 21, 15]. To our knowledge, however, a self-contained account that combines exact balance laws, a retained-spin Chapman–Enskog construction, an irreducible first-order sector decomposition, and explicit dilute-gas rough-sphere estimates for the rotational coefficients has not been readily available in one place. The role of the present paper is therefore primarily reconstructive and expository: it collects these strands into a single notation and makes explicit which parts of the retained-spin closure are exact, which are formal Chapman–Enskog results, and which remain estimate-level rough-sphere evaluations.
The aim of the present paper is therefore narrow but useful. We do not attempt to develop a new response theory or a new application. The numerical material added here remains deliberately targeted rather than exhaustive: expanded EDMD benchmarks are used only to test the dilute rough-sphere estimates rather than to launch a separate transport-simulation program. We otherwise write out a detailed derivation of the retained-spin closure from the Boltzmann–Curtiss equation, with special attention to the exact balance laws, the irreducible tensor decomposition of the first-order source, the formal first-order coefficient problem, and the dilute-gas estimates of the key rotational coefficients for perfectly rough elastic hard spheres. The paper is intended to be readable line-by-line by a reader who wants to check the intermediate manipulations. A companion manuscript Ref. [25] addresses the response-theoretic consequences of the retained-spin closure, including EDMD observability and model discrimination among retained-spin, eliminated-spin, and polynomial higher-gradient descriptions. The present paper is instead derivation- and coefficient-estimate-oriented: its purpose is to provide the exact balance laws, the generalized Chapman–Enskog construction with retained spin, and dilute-gas rough-sphere estimates for selected key rotational coefficients.
@@
    We derive the exact mass, momentum, and intrinsic-spin balances from the Boltzmann–Curtiss equation, with the stress and couple stress defined as moments of the distribution function.
@@
Two limitations should be stated at the outset. First, the present Chapman–Enskog construction is generalized or extended: the mean spin is treated as quasi-slow rather than as a strict collision invariant. More precisely, the retained-spin ordering adopted below assumes that the residual spin-relaxation source of the local quasi-equilibrium manifold is $O(\varepsilon)$, i.e. of the same asymptotic order as the first gradient corrections. Second, the antisymmetric stress channel is structurally closed at first order, but a fully coefficient-complete evaluation of $\eta_{r}$ for a rough-sphere collision operator would require an explicit collisional-transfer bracket. The rough-sphere formulas derived here should therefore be read as controlled low-density estimates rather than as the final word on the full antisymmetric Chapman–Enskog problem.
@@ II.3 Collision invariants
The total-energy invariant will not be used explicitly below and is therefore omitted from the displayed list. Equations (8)–(10) are the only collision properties needed for the exact balance derivation.
@@ III.2 Linear-momentum balance
For structured or finite-size particles, a coarse-grained continuum momentum balance may contain an additional collisional-transfer contribution $\sigma_{ij}^{(\mathrm{int})}$ beyond the one-particle kinetic moment. It is useful to reserve the notation
@@
for the total Cauchy stress used later in the constitutive closure. In a point-particle theory $\sigma_{ij}^{(\mathrm{int})}\equiv 0$.
@@ III.3 Total and intrinsic angular momentum
Likewise,
$\mathcal{J}_{ij}$ $=\varepsilon_{ipq}x_{p}\Pi^{(k)}_{qj}+\int(\mathbb{I}\bm{\omega})_{i}v_{j}f\,\mathrm{d}\Gamma$
$=\varepsilon_{ipq}x_{p}\Pi^{(k)}_{qj}+\rho J\omega_{0i}u_{j}+m_{ij},$ (46)
where the couple stress is defined by the nonconvective intrinsic flux
$m_{ij}:=\int(\mathbb{I}\bm{\omega})_{i}(v_{j}-u_{j})f\,\mathrm{d}\Gamma=\int(\mathbb{I}\bm{\omega})_{i}c_{j}f\,\mathrm{d}\Gamma.$ (47)
Next derive the balance of orbital angular momentum directly from the linear momentum equation. Starting from Eq. (26), multiply by $\varepsilon_{ipq}x_{p}$:
$\partial_{t}(\varepsilon_{ipq}x_{p}\rho u_{q})+\varepsilon_{ipq}x_{p}\partial_{x_{j}}\Pi^{(k)}_{qj}$ $=\rho(\bm{x}\times\bm{F})_{i}.$ (48)
@@
$\varepsilon_{ipq}x_{p}\partial_{x_{j}}\Pi^{(k)}_{qj}=\partial_{x_{j}}(\varepsilon_{ipq}x_{p}\Pi^{(k)}_{qj})-\varepsilon_{ijq}\Pi^{(k)}_{qj}$, (49)
@@
$\partial_{t}(\varepsilon_{ipq}x_{p}\rho u_{q})+\partial_{x_{j}}(\varepsilon_{ipq}x_{p}\Pi^{(k)}_{qj})=\rho(\bm{x}\times\bm{F})_{i}+\varepsilon_{ijq}\Pi^{(k)}_{qj}$. (50)
Subtracting Eq. (50) from Eq. (42) and using Eqs. (45)–(46) gives the exact intrinsic-spin balance
$\partial_{t}(\rho J\omega_{0i})+\partial_{x_{j}}(\rho J\omega_{0i}u_{j}+m_{ij})=n\tau_{i}^{\mathrm{ext}}.$ (51)
Defining the body couple per unit mass by
$\rho G_{i}:=n\tau_{i}^{\mathrm{ext}},$ (52)
we may write
$\partial_{t}(\rho J\bm{\omega}_{0})+\bm{\nabla}\!\cdot(\rho J\bm{\omega}_{0}\otimes\bm{u}+\bm{m})=\rho\bm{G}.$ (53)
@@
$\rho J\frac{\mathrm{D}\omega_{0i}}{\mathrm{D}t}=\partial_{x_{j}}m_{ij}+\rho G_{i}.$ (54)
@@ III.4 Where does the antisymmetric stress appear?
Equation (54) is an exact moment identity written with the nonconvective intrinsic flux $m_{ij}$. No explicit term of the form $\varepsilon_{ijk}\sigma_{jk}$ appears because the intrinsic balance was obtained by subtracting the orbital balance from the total-angular-momentum balance. In that subtraction the orbital–intrinsic exchange is already accounted for.
In micropolar continuum mechanics it is often convenient to rewrite the closed spin balance as
$\rho J\frac{\mathrm{D}\omega_{0i}}{\mathrm{D}t}=\partial_{x_{j}}m_{ij}+\tau_{i}^{(\sigma)}+\rho G_{i},\qquad\tau_{i}^{(\sigma)}:=-\varepsilon_{ijk}\sigma_{jk}.$ (55)
This is a constitutive rewriting of the same exchange, not an additional independent source. One may represent the orbital–intrinsic transfer either through the antisymmetric part of the total stress or through an equivalent spin-torque term, but not both independently. We will return to this point in Sec. V.
@@ IV.2 Local quasi-equilibrium distribution
$m_{ij}^{(0)}=\int I\omega_{i}c_{j}f^{(0)}\,\mathrm{d}\Gamma=0,$ (61)
@@
At this point the extended-hydrodynamic ordering should be stated explicitly. For a strict classical Chapman–Enskog construction, the local Maxwellian lies on the exact null manifold of the collision operator, so that $C[f^{(0)},f^{(0)}]=0$. For the retained-spin reference state (59), however, the noninvariant spin channel need not vanish identically. We therefore introduce an auxiliary bookkeeping parameter $\delta_{s}$ that measures the residual spin-relaxation strength on the retained-spin manifold and adopt the matched ordering
$\delta_{s}=O(\varepsilon),\qquad C[f^{(0)},f^{(0)}]=\delta_{s}\,\mathcal{R}^{(\mathrm{ax})}[f^{(0)}].$
Here $\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]$ denotes the residual axial part of the collision operator on the retained-spin manifold. It is orthogonal to the strict one-particle collision invariants,
$\int\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]\,\mathrm{d}\Gamma=0,\qquad\int mv_{i}\,\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]\,\mathrm{d}\Gamma=0,\qquad\int\left(\frac{m}{2}v^{2}+\frac{I}{2}\omega^{2}\right)\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]\,\mathrm{d}\Gamma=0,$
but its intrinsic-spin moment need not vanish. Operationally, the present retained-spin closure may therefore be regarded as a double expansion in $\varepsilon$ and $\delta_{s}$, with the bookkeeping $\delta_{s}\sim\varepsilon$. Under this matched ordering, the residual spin-relaxation source enters at the same asymptotic order as the usual gradient-driven first-order source. This is the precise sense in which $\bm{\omega}_{0}$ is treated as a quasi-slow variable in the present paper. If instead $\delta_{s}=O(1)$, the irreducible sector decomposition derived below still identifies the correct constitutive channels, but the construction should then be read as a formal extended-hydrodynamic closure rather than as a strict one-parameter Chapman–Enskog limit. The explicit rough-sphere estimates derived later are therefore coefficient evaluations within the same retained-spin constitutive structure; they are not, by themselves, an additional proof of the ordering $\delta_{s}\sim\varepsilon$.

### 13 Apr 2026 v2

@@ Abstract
We derive a retained-spin micropolar hydrodynamic closure from the Boltzmann–Curtiss equation using a generalized Chapman–Enskog construction in which the local mean spin is retained as a quasi-slow variable. Starting from the one-particle kinetic balance identities and the corresponding exact coarse-grained finite-size balances for mass, linear momentum, and intrinsic angular momentum, we keep the collisional-transfer contribution to the antisymmetric stress explicit in the spin balance, decompose the first-order source into irreducible scalar, axial, and symmetric-traceless sectors, and show explicitly how the standard micropolar constitutive structure with coefficients $(\eta,\xi,\eta_{r},\alpha,\beta,\gamma)$ emerges. This decomposition makes clear that the one-particle kinetic stress contributes only to the symmetric stress, whereas the rotational viscosity belongs to a collisional-transfer channel. For perfectly rough elastic hard spheres, we further obtain explicit dilute-gas estimates for the rotational viscosity $\eta_{r}$ from homogeneous spin relaxation and for the transverse spin-diffusion combination $\beta+\gamma$ from a transport-relaxation calculation. Targeted event-driven molecular-dynamics simulations are used as a posteriori checks: expanded homogeneous-spin density and roughness sweeps support the predicted $n^{2}$ and $K/(K+1)$ trends for $\eta_{r}$, while finite-$k$ transverse runs provide a qualitative diagnostic of the retained-spin response. The result is a self-contained derivation and coefficient-level estimate of retained-spin micropolar hydrodynamics that clarifies which parts of the closure are exact balance-law statements, which are first-order generalized Chapman–Enskog results, and which remain controlled rough-sphere estimates.
@@ I Introduction
The basic links between the kinetic and continuum descriptions are classical, but in practice the derivation is scattered across several literatures. The exact moment balances are often stated without algebra. The Chapman–Enskog construction is typically described only in outline. The antisymmetric stress channel, which is the key to rotational viscosity, is especially susceptible to confusion because it is not generated by the symmetric one-particle kinetic stress and, for finite-size rough particles, must already be kept explicit in the exact coarse-grained intrinsic-spin balance. In addition, when a local spin field is retained at the hydrodynamic level, the calculation is no longer a strict hydrodynamic reduction in the narrow sense: the mean spin is not a collision invariant in general and must be treated as a quasi-slow variable in the sense of extended thermodynamics [19, 13].
@@
Subsequent developments also proceeded largely along separate tracks. Micropolar continuum theory matured as an independent constitutive framework [10, 11]; granular and rough-sphere kinetic theories emphasized rapid granular flow, rough-particle transport, and hydrodynamic closures for particles with rotational degrees of freedom [14, 12, 17]; and recent Boltzmann–Curtiss-based studies have focused either on local-spin first-order closures or on higher-order constitutive models for polyatomic gases [26, 21, 15]. To our knowledge, however, a self-contained account that combines the exact balance-law structure, a retained-spin Chapman–Enskog construction, an irreducible first-order sector decomposition, and explicit dilute-gas rough-sphere estimates for the rotational coefficients has not been readily available in one place. The role of the present paper is therefore primarily reconstructive and expository: it collects these strands into a single notation and makes explicit which parts of the retained-spin closure are exact, which are formal Chapman–Enskog results, and which remain estimate-level rough-sphere evaluations.
The aim of the present paper is therefore narrow but useful. We do not attempt to develop a new response theory or a new application. The numerical material added here remains deliberately targeted rather than exhaustive: expanded EDMD benchmarks are used only to test the dilute rough-sphere estimates rather than to launch a separate transport-simulation program. We otherwise write out a detailed derivation of the retained-spin closure from the Boltzmann–Curtiss equation, with special attention to the exact balance-law structure, the irreducible tensor decomposition of the first-order source, the formal first-order coefficient problem, and the dilute-gas estimates of the key rotational coefficients for perfectly rough elastic hard spheres. The paper is intended to be readable line-by-line by a reader who wants to check the intermediate manipulations. A companion manuscript Ref. [25] addresses the response-theoretic consequences of the retained-spin closure, including EDMD observability and model discrimination among retained-spin, eliminated-spin, and polynomial higher-gradient descriptions. The present paper is instead derivation- and coefficient-estimate-oriented: its purpose is to provide the exact balance-law structure with the collisional-transfer torque channel made explicit, the generalized Chapman–Enskog construction with retained spin, and dilute-gas rough-sphere estimates for selected key rotational coefficients.
@@
    We derive the one-particle kinetic balance identities from the Boltzmann–Curtiss equation and write the corresponding exact finite-size intrinsic-spin balance with the collisional-transfer stress torque made explicit.
@@
Two limitations should be stated at the outset. First, the present Chapman–Enskog construction is generalized or extended: the mean spin is treated as quasi-slow rather than as a strict collision invariant. More precisely, the retained-spin ordering adopted below assumes that the residual axial relaxation of the retained-spin manifold, represented in the collisional-transfer channel, is $O(\varepsilon)$, i.e. of the same asymptotic order as the first gradient corrections. Second, the antisymmetric stress channel is structurally closed at first order, but a fully coefficient-complete evaluation of $\eta_{r}$ for a rough-sphere collision operator would require an explicit collisional-transfer bracket. The rough-sphere formulas derived here should therefore be read as controlled low-density estimates rather than as the final word on the full antisymmetric Chapman–Enskog problem.
@@ II.3 Collision invariants
The total-energy invariant will not be used explicitly below and is therefore omitted from the displayed list. Equations (8)–(10) are sufficient for the one-particle moment identities derived below. For finite-size rough particles, however, the exact local coarse-grained momentum and angular-momentum balances contain additional contact-transfer fluxes. In particular, the antisymmetric part of the total stress produces a local intrinsic-spin torque. We therefore distinguish below between the one-particle kinetic moments obtained directly from Eq. (2) and the exact finite-size balance laws used for the retained-spin micropolar closure.
@@ III.2 Linear-momentum balance
For structured or finite-size particles, Eq. (32) identifies only the one-particle kinetic contribution. The exact coarse-grained local momentum balance contains an additional contact-transfer stress $\sigma_{ij}^{(\mathrm{int})}$. We therefore reserve the notation
@@
for the total Cauchy stress and define the corresponding total momentum flux by
$\Pi_{ij}:=\rho u_{i}u_{j}-\sigma_{ij}=\Pi_{ij}^{(k)}-\sigma_{ij}^{(\mathrm{int})}.$ (34)
The exact finite-size momentum balance is then
$\partial_{t}(\rho u_{i})+\partial_{x_{j}}\Pi_{ij}=\rho F_{i},\qquad\rho\frac{\mathrm{D}u_{i}}{\mathrm{D}t}=\partial_{x_{j}}\sigma_{ij}+\rho F_{i}.$ (35)
In a point-particle theory $\sigma_{ij}^{(\mathrm{int})}\equiv 0$ and Eq. (35) reduces to Eq. (32).
@@ III.3 Total and intrinsic angular momentum
Likewise, the one-particle part of the angular-momentum flux is
$\mathcal{J}^{(k)}_{ij}$ $=\varepsilon_{ipq}x_{p}\Pi^{(k)}_{qj}+\int(\mathbb{I}\bm{\omega})_{i}v_{j}f\,\mathrm{d}\Gamma$
$=\varepsilon_{ipq}x_{p}\Pi^{(k)}_{qj}+\rho J\omega_{0i}u_{j}-m_{ij}^{(k)},$ (48)
where the kinetic couple stress is defined as minus the nonconvective intrinsic flux
$m_{ij}^{(k)}:=-\int(\mathbb{I}\bm{\omega})_{i}(v_{j}-u_{j})f\,\mathrm{d}\Gamma=-\int(\mathbb{I}\bm{\omega})_{i}c_{j}f\,\mathrm{d}\Gamma.$ (49)
For finite-size particles the exact coarse-grained total angular-momentum flux contains additional contact-transfer contributions. We therefore write
$\mathcal{J}_{ij}:=\varepsilon_{ipq}x_{p}\Pi_{qj}+\rho J\omega_{0i}u_{j}-m_{ij},$ (50)
where $m_{ij}$ is the exact intrinsic flux. In the dilute first-order closure developed below, its leading contribution is represented by the one-particle moment $m_{ij}^{(k)}$.
Next derive the balance of orbital angular momentum directly from the exact finite-size momentum equation (35). Multiplying by $\varepsilon_{ipq}x_{p}$ gives
$\partial_{t}(\varepsilon_{ipq}x_{p}\rho u_{q})+\varepsilon_{ipq}x_{p}\partial_{x_{j}}\Pi_{qj}$ $=\rho(\bm{x}\times\bm{F})_{i}.$ (51)
@@
$\varepsilon_{ipq}x_{p}\partial_{x_{j}}\Pi_{qj}=\partial_{x_{j}}(\varepsilon_{ipq}x_{p}\Pi_{qj})-\varepsilon_{ijq}\Pi_{qj},$ (52)
@@
$\partial_{t}(\varepsilon_{ipq}x_{p}\rho u_{q})+\partial_{x_{j}}(\varepsilon_{ipq}x_{p}\Pi_{qj})=\rho(\bm{x}\times\bm{F})_{i}+\varepsilon_{ijq}\Pi_{qj}.$ (53)
The corresponding exact coarse-grained balance of total angular momentum reads
$\partial_{t}\mathcal{L}_{i}+\partial_{x_{j}}\mathcal{J}_{ij}=\rho(\bm{x}\times\bm{F})_{i}+\rho G_{i},$ (54)
where, as before,
$\rho G_{i}:=n\tau_{i}^{\mathrm{ext}}.$ (55)
Subtracting Eq. (53) from Eq. (54) and using Eqs. (47) and (50) gives the exact finite-size intrinsic-spin balance
$\partial_{t}(\rho J\omega_{0i})+\partial_{x_{j}}(\rho J\omega_{0i}u_{j}-m_{ij})=-\varepsilon_{ijq}\Pi_{qj}+\rho G_{i}.$ (56)
Since the convective part $\rho u_{q}u_{j}$ is symmetric, $\varepsilon_{ijq}\Pi_{qj}=-\varepsilon_{ijk}\sigma_{jk}$, and we may write
$\partial_{t}(\rho J\bm{\omega}_{0})+\bm{\nabla}\!\cdot(\rho J\bm{\omega}_{0}\otimes\bm{u}-\bm{m})=\bm{\tau}^{(\sigma)}+\rho\bm{G},\qquad\tau_{i}^{(\sigma)}:=-\varepsilon_{ijk}\sigma_{jk}.$ (57)
@@
$\rho J\frac{\mathrm{D}\omega_{0i}}{\mathrm{D}t}=\partial_{x_{j}}m_{ij}-\varepsilon_{ijk}\sigma_{jk}+\rho G_{i}.$ (58)
@@ III.4 Where does the antisymmetric stress appear?
Equation (58) already contains the antisymmetric-stress torque through the exact total stress. In a point-particle theory $\sigma_{ij}=\sigma_{ij}^{(k)}$ is symmetric, so $\tau_{i}^{(\sigma)}=0$ and the intrinsic balance reduces to the purely kinetic flux form. For finite-size rough particles, however, the contact-transfer part $\sigma_{ij}^{(\mathrm{int})}$ need not be symmetric, and
$\tau_{i}^{(\sigma)}:=-\varepsilon_{ijk}\sigma_{jk}=-\varepsilon_{ijk}\sigma_{jk}^{(\mathrm{int})}$ (59)
represents the local orbital–intrinsic exchange. This is the exact channel that later carries the rotational viscosity $\eta_{r}$; it is not an additional constitutive term to be appended after the fact. We return to this point in Sec. V.
@@ IV.2 Local quasi-equilibrium distribution
$m_{ij}^{(k,0)}=-\int I\omega_{i}c_{j}f^{(0)}\,\mathrm{d}\Gamma=0,$ (65)
@@
At this point the extended-hydrodynamic ordering should be stated explicitly. For a strict classical Chapman–Enskog construction, the local Maxwellian lies on the exact null manifold of the one-particle collision operator, so that $C[f^{(0)},f^{(0)}]=0$. In the retained-spin closure for finite-size rough particles, however, the axial collisional-transfer channel associated with the exact spin torque need not vanish identically on the retained-spin manifold. We therefore introduce an auxiliary bookkeeping parameter $\delta_{s}$ that measures the residual axial relaxation strength and write the matched-order axial source symbolically as
$\delta_{s}=O(\varepsilon),\qquad\mathcal{S}^{(\mathrm{ax})}_{i}a_{i}=\mathcal{S}^{(\mathrm{ax},\nabla)}_{i}a_{i}+\delta_{s}\,\mathcal{R}^{(\mathrm{ax})}[f^{(0)}].$
Here $\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]$ denotes the residual axial contribution of the retained-spin manifold, understood to belong to the collisional-transfer channel rather than to the symmetric one-particle kinetic stress. It is orthogonal to the strict mass, linear-momentum, and energy invariants,
$\int\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]\,\mathrm{d}\Gamma=0,\qquad\int mv_{i}\,\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]\,\mathrm{d}\Gamma=0,\qquad\int\left(\frac{m}{2}v^{2}+\frac{I}{2}\omega^{2}\right)\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]\,\mathrm{d}\Gamma=0.$
Operationally, the present retained-spin closure may therefore be regarded as a double expansion in $\varepsilon$ and $\delta_{s}$, with the bookkeeping $\delta_{s}\sim\varepsilon$. Under this matched ordering, the residual axial relaxation enters at the same asymptotic order as the usual gradient-driven first-order source. This is the precise sense in which $\bm{\omega}_{0}$ is treated as a quasi-slow variable in the present paper. If instead $\delta_{s}=O(1)$, the irreducible sector decomposition derived below still identifies the correct constitutive channels, but the construction should then be read as a formal extended-hydrodynamic closure rather than as a strict one-parameter Chapman–Enskog limit. The explicit rough-sphere estimates derived later are therefore coefficient evaluations within the same retained-spin constitutive structure; they are not, by themselves, an additional proof of the ordering $\delta_{s}\sim\varepsilon$.